### v1

# Feasibility of Federated Learning from Client Databases with Different Brain Diseases and MRI Modalities

Abstract
Segmentation models for brain lesions in MRI are commonly developed for a specific disease and trained on data with a predefined set of MRI modalities. Each such model cannot segment the disease using data with a different set of MRI modalities, nor can it segment any other type of disease. Moreover, this training paradigm does not allow a model to benefit from learning from heterogeneous databases that may contain scans and segmentation labels for different types of brain pathologies and diverse sets of MRI modalities. Is it feasible to use Federated Learning (FL) for training a single model on client databases that contain scans and labels of different brain pathologies and diverse sets of MRI modalities? We demonstrate promising results by combining appropriate, simple, and practical modifications to the model and training strategy: Designing a model with input channels that cover the whole set of modalities available across clients, training with random modality drop, and exploring the effects of feature normalization methods.  Evaluation on 7 brain MRI databases with 5 different diseases shows that such FL framework can train a single model that is shown to be very promising in segmenting all disease types seen during training. Importantly, it is able to segment these diseases in new databases that contain sets of modalities different from those in training clients.  These results demonstrate, for the first time, feasibility and effectiveness of using FL to train a single segmentation model on decentralised data with diverse brain diseases and MRI modalities, a necessary step towards leveraging heterogeneous real-world databases. Code will be made available at: https://github.com/FelixWag/FL-MultiDisease-MRI
Keywords:
multi-modal MRI brain lesions federated learning
1 Introduction
Brain MRI segmentation is crucial for disease analysis, diagnosis and treatment. Traditionally, brain MRI segmentation models are trained on data for a specific disease, acquired with a predefined protocol and set of MRI modalities [8, 9, 30, 28]. Models trained with this standard approach cannot segment other types of diseases, and expect the same modalities in all training and testing data. While a recent study [26] demonstrated the feasibility of training a segmentation model across multiple brain MRI databases with different diseases and sets of modalities, that study used a centralised database. It did not address the practical challenge related to the sensitive nature of medical data, which prevents data sharing between institutions and therefore the creation of an appropriately diverse centralised database due to legal and privacy concerns.
To address this, this study aims to develop a Federated Learning (FL) framework [15] for training a single model on diverse brain lesion databases, each with a different disease and set of MRI modalities. Central questions for this study are: How to make such FL training feasible? Will it be beneficial compared to training a disease-specific model? Will it result in a model that can segment new data that have different combination of modalities than the training data?
Related Work: While previous studies investigated training with heterogeneous databases [27, 22, 16], they have not studied FL using decentralised databases, nor data with diverse sets of MRI modalities. A recent study [26] investigated heterogeneous databases with diverse sets of MRI modalities in a centralized scenario. In contrast, our study focuses on the challenging setting of decentralized FL, which has not been explored yet. Some works have focused on FL of a segmentation model for a single, specific disease when a specific set of multi-modal MRI is available in each client [11, 19, 17]. Other works studied FL for learning from data with different labels in segmentation of CT [20] or classification in MRI [6]. In summary, there are works that focus on FL with multi-modal MRI and FL for heterogeneous tasks; however, the combination of both, especially with different sets of modalities per client, has never been explored.
Learning from databases with varying sets of MRI modalities relates to research on developing multi-modal models that perform well even when a modality is missing at inference time [4, 18, 21, 5, 31, 24]. This line of work focuses on learning a model for a specific disease, from data with the same set of modalities during training. While addressing missing modalities is an important capability of our method, it is not the primary topic of this research. Instead, this study concentrates on the effectiveness and practicality of learning from diverse databases with different sets of input modalities and segmentation tasks.
Contribution: This study develops a simple but practical solution for training via FL a single model that can process different sets of modalities and segment brain lesions of different types. We show this can be achieved by a standard U-net, enhanced such that input channels can receive the whole set of unique modalities available across clients and using random drop of modalities during training to prevent associations of modalities with specific pathologies. Additionally, due to the high data distribution shift across diverse databases, we present an analysis of the effects of feature normalization techniques. Experimentation on 7 brain MRI databases with 5 different pathologies and 6 MRI modalities provides the following findings:
  • Feasibility: This study is the first to show feasibility of Federated training of a single segmentation model on decentralised databases with different brain diseases and sets of MRI modalities.
  • Knowledge Transfer Across Pathologies: This paper demonstrates effectiveness of FL for jointly learning to segment multiple pathologies, achieving performance that compares favorably to disease-specific models.
  • Segmenting new combination of modalities: This study shows that models trained with this FL paradigm can segment data from new databases with combination of modalities not seen during training.
Figure 1: Method overview: (1) Initialization of model with distinct set of modalities across clients. (2) Federated training of the model with random modality drop. (3) Generalization to unseen clients with different input MRI modality combination.
2 Method
2.1 Problem setting: Multi-Modal MRI FL for Multiple Tasks:
We consider $C$ multi-modal clients, where each client $c$ ($1\leq c\leq C$) is associated with a set of MRI-modalities $\mathcal{M}_{c}$ and a dataset $\mathcal{D}_{c}=\{(\mathcal{X}_{ci},y_{ci})\}_{i=1}^{|\mathcal{D}_{c}|}$, where $\mathcal{X}_{ci}=\{{x}_{ci}^{m}\}_{m=1}^{|\mathcal{M}_{c}|}$ denotes the multi-modal MRI image and $y_{ci}$ the corresponding segmentation mask of the $i$-th sample. Here, each $x_{ci}^{m}\in\mathbb{R}^{w\times h\times d}$, represents a 3D image. Furthermore, $\mathcal{M}=\bigcup_{i=1}^{C}M_{i}$ denotes the set that represents all unique modalities across all clients and $p:=|\mathcal{M}|$ is the total number of unique modalities across all datasets. Each client solves a specific segmentation task $\mathcal{T}_{c}$. Our goal is train, without data sharing, a single global model parameterized by $\theta$, that performs well on all tasks and clients. The global objective is to minimize the sum of the client-local loss functions:
$\min_{\theta}[\mathcal{L}(\theta):=\frac{1}{C}\sum_{c=1}^{C}\mathcal{L}_{c}(\theta;\mathcal{D}_{c};\mathcal{T}_{c})],$ (1)
where $\mathcal{L}_{c}$ denotes the local loss function of client $c$. As loss functions, we use a weighted combination of Dice and pixel-wise Binary Cross-Entropy (BCE) losses $\mathcal{L}_{c}=\alpha\mathcal{L}_{Dice}+(1-\alpha)\mathcal{L}_{BCE}$, where $\alpha=0.8$ in our experiments. In the standard FL setting with $E$ communication rounds, a global server initializes a global model and communicates it to every client. In each round, each client performs $\tau$ iterations of updates on its local data and loss function $\mathcal{L}_{c}(\theta;\mathcal{D}_{c};\mathcal{T}_{c})$. Afterwards, every client sends their updated model back to the server. The server combines these updates, typically by averaging (FedAvg), into a new global model before starting a new communication round [15]. This allows decentralised training without data sharing. We want to highlight the differences from the standard FL setting: this setting varies not only in the input space $|\mathcal{M}|>1$, indicating multiple modalities, but also in the label space $|\bigcup_{i=1}^{C}T_{i}|>1$, involving segmenting different brain pathologies per client. Additionally, this setting considers different sets of modalities per client, an aspect mostly overlooked in FL.
2.2 Training a Unified Model for Multiple Segmentation Tasks
Handling Different Sets of Modalities Across Clients: Since every client comes with a different combination of MRI modalities, the challenge arises to find a single model architecture capable of processing different sets of input modalities. To address this, we augment each client’s input to cover the full set of distinct MRI modalities across all clients, and fill missing modalities $m\in\mathcal{M}-\mathcal{M}_{c}$ of each client $c$ with zeros. This transforms the input space of each client to $\mathcal{X}_{c}\in\mathbb{R}^{p\times w\times h\times d}$. Consequently, our global model accepts the set of distinct MRI modalities as input and becomes a function $f(\cdot;\theta):\mathbb{R}^{p\times w\times h\times d}\to\mathbb{R}^{w\times h\times d}$, capable of processing each client’s input modality combination, ensuring compatibility across all clients.
Modality Drop for Pathology Disassociation & Missing Modalities: To develop a model that generalises to modality combinations not seen during training, we introduce modality drop, where we randomly set input modalities to zero during training. Therefore, we define a probabilistic masking function $Z(\mathcal{X},\phi)$ where each input modality $x^{m}\in\mathcal{X}$ is set to zero with probability $\phi$. This prevents the model from associating specific modality combinations with specific pathologies, ensuring it can accurately segment known pathologies in unseen clients with modality combinations not seen during training. Furthermore, it also enhances the model’s robustness against missing modalities during testing, as shown in our evaluation.
Feature Normalization for Heterogeneous Clients: Each client’s data contains different pathologies and modalities, which increases the difficulty of training a model due to data heterogeneity. To tackle this, we employ Batch Normalization (BN) [7], but keep client-specific BN parameters and statistics, and exclude them from the model averaging step, as proposed in [1, 12]. Our experiments show that this method proves effective in improving the performance of the training (source) databases. However, it is not clear how to apply the source-specific BN layer parameters/statistics to unseen (target) clients during inference. One way involves averaging the learned BN parameters and estimating running statistics on the target data, which has the disadvantage that target data is needed. Therefore, in our experiments, we also explore alternative feature normalization methods for this training framework with inherently high data heterogeneity. Specifically, the recently proposed Normalization Free (NF) technique [32] uses re-parameterized convolutional layers instead of explicitly normalizing statistics of feature activations. This does not require target data to estimate the target statistics, alleviating this requirement of keeping client-specific BN layers. Empirical analysis (Sec. 3) also shows that this results in methods that generalize better to unseen clients. However, it slightly decreases performance on source clients. Therefore, the choice of feature normalization method in this framework should be chosen based on the target application for the trained model (generalization to new domains or application to source domains) and whether data are expected to have been collected a-priori from the target domain for estimating BN statistics.
Algorithm: To summarise, our simple yet effective algorithm is shown in Fig. 1. Initially, the central server receives a list of available MRI modalities from each client’s dataset. Afterwards, the server initializes a model architecture that accepts the set of unique MRI modalities $\mathcal{M}$ as input and the FL algorithm starts (Sec. 2.1). During training, each client’s input modalities are augmented to the whole set of unique MRI modalities and modality drop is applied.
3 Experiments & Results
3.1 Datasets & Prepocessing
We use 7 annotated brain MRI databases, each with different MRI modalities and segmentation tasks collected from different sites: ATLAS v2.0 [13] for stroke lesion, BRATS 2016 [2] for tumor, MSSEG [3] for multiple sclerosis, TBI from our site for traumatic brain injury, WMH [10] for White Matter Hyperintensity, ISLES 2015’s SISS subtask [14] for stroke lesion, and Tumor2 for tumour segmentation from our institution. The data is split into training and validation data. Tab. 1 gives an overview of the data sizes and modalities. We resample the images to 1x1x1 mm, skull-strip and apply z-score intensity normalization, and combine all lesion types across databases into a single ’lesion’ label.
Table 1: MRI databases overview: ISLES and Tumor2 are used for evaluation only, except for single center training, where a 70-30 train-validation split is used.
Database Modalities Disease/Task #Samples #Val. Samples
T1 T1c FLAIR T2 PD SWI
BRATS Tumor 484 40
MSSEG Multiple Sclerosis 53 16
ATLAS Stroke Lesion 654 195
TBI Traumatic Brain Injury 281 125
WMH White Matter Hyperintensity 60 18
ISLES Stroke Lesion 28 28
Tumor2 Tumor 57 57
3.2 Experimental Setup
Setup & Methods: We use the Residual U-Net [29] as network architecture and Adam optimizer. We linearly scale the learning rate from 0 to $10^{-3}$ over 50 communication rounds, keep it constant until communication round 150, and then linearly decay it. We set $\tau=58$ (equivalent to 1 epoch for the largest database, ATLAS), $E=300$ (or 300 epochs for centralised training), and use a batch size of 8. The Dice score on the validation set is reported after the last epoch. For our federated experiments, both the standard FedAvg algorithm [15] and, separately, FedBN [12] (to keep client-specific BN layers) are used. As discussed in Section 2.2, in addition to BN, our experiments investigate the impact of alternative feature normalization methods that do not rely on client-specific normalization layers: Group Norm (GN) [25], Instance Norm (IN) [23] and the NF [32] technique.
3.3 Results & Discussion
Unified Model - Improves Performance Across Databases: Results in Tab. 2 show that, when keeping client-specific BN parameters, federated joint training across diverse MRI databases with different brain diseases and different input modalities improves the model’s performance on each database, either surpassing or matching single center training. This highlights the benefits of transferring knowledge between different brain diseases, especially for databases with limited data (MSSEG, TBI, and WMH). Training with modality drop and client-specific BN layers leads to the highest overall results and improves over FedAvg using averaged BN parameters, demonstrating the benefits of client specific BN layers for source clients. Furthermore, this joint learning framework produces a single model instead of training 5 individual models, demonstrating its practicality.
Table 2: Separate models are trained and evaluated on each database (single center), with only BN results shown as they yield the best results. In the federated setting the model is trained across ATLAS, BRATS, MSSEG, TBI, and WMH with various feature normalization methods. Furthermore, results of modality drop are reported.
Setting Mod.Drop Norm ATLAS BRATS MSSEG TBI WMH Average
Single Database BN 52.8 91.9 68.3 56.1 71.5 68.1
Federated (client spec. BN params) BN 50.9 91.8 70.2 55.8 74.8 68.7
Single Database BN 52.8 91.7 66.7 53.9 69.1 66.8
Federated (avg. BN params) BN 43.1 89.4 63.2 51.4 67.3 62.9
Federated IN 49.0 90.7 66.6 52.9 71.0 66.0
Federated GN 48.1 91.0 67.3 53.7 70.4 66.1
Federated NF 51.2 91.1 64.8 54.4 72.3 66.8
Federated (client spec. BN params) BN 54.5 91.8 69.1 56.2 73.7 69.1
Modality Drop Improves Robustness Against Missing Modalities: Our initial step towards developing a model capable of segmenting pathologies with modality combinations different from those seen during training involves robustness against missing modalities. Results in Tab. 3 show the need for modality drop to improve the model’s robustness against missing modalities. When using modality drop during training, the Dice score decreases by only 3.9 when evaluating with missing modalities compared to all modalities, versus a 21.3 decrease without modality drop. Without modality drop, the model relies on having all modalities for a specific disease present during inference. Therefore, the following experiments focus on models trained with modality drop.
Table 3: Results from jointly training a model using client-specific BN layers on ATLAS, BRATS, MSSEG, TBI, and WMH, both with and without modality drop and evaluated on the same databases by randomly excluding modalities during testing to mimic missing modalities, as well as with all modalities present during testing. Training with modality drop is important so that at test-time the model can work with different combinations of modalities. The average Dice score across all databases is reported.
Training Method Evaluation (Dice) $\Delta$ Dice
Modality Drop No Modality Drop
No Modality Drop 47.4 68.7 21.3
Modality Drop 65.2 69.1 03.9
Zero-Shot Generalization to Unseen Modality Combinations: Tab. 4 presents generalization results on unseen databases with modality combinations not seen during training. It is remarkable that despite not seeing any of the new client’s data, the best federated model’s average Dice score reaches only 4.1% below the upper bound of single center training on the target database and comes close to centralized joint training (for which only results with IN are shown as it achieved best performance). Traditionally, models are trained on databases specific to a disease and evaluated on unseen clients with the same disease, as we trained with ATLAS (for stroke lesion) and evaluated on ISLES, and trained with BRATS (for brain tumors) and evaluated on Tumor2, setting this as baseline. Both are only trained and evaluated on T1 since it is the only overlapping modality. Federated joint training achieves a performance gain of 49.3 on ISLES and 8.1 on Tumor2, reinforcing the benefits of joint training across multiple MRI databases compared to pathology specific models.
Additionally, modality drop proves essential to dissociate specific modality combinations with certain pathologies, as all jointly trained models with modality drop show improved average performance on segmenting pathologies with modality combinations not seen during training. For Tumor2, there is extensive improvement (both in centralised and federated settings), whereas for ISLES, the federated models performed unexpectedly slightly worse, likely due to gradient stochasticity in federated optimization.
Comparing client-specific BN layers to non-client-specific normalization approaches, the results indicate that the NF technique outperforms all other techniques. Using the average of client-specific BN layers and estimating the running statistics on the validation data gives the second-best results. However, this estimation is less practical as it requires unlabeled data from the unseen client.
Table 4: Generalization results on unseen clients ISLES and Tumor2, each with modality combinations not seen during training (reported in Dice score): Separate models were trained on ISLES and Tumor2 and evaluated on the same database (Upper Bound). Baselines involve training on stroke lesion database ATLAS, evaluating on ISLES, and training on tumor database BRATS, evaluating on Tumor2. These results are compared to FL, using different normalization techniques, across 5 training databases (ATLAS, BRATS, MSSEG, TBI, and WMH) and evaluated on ISLES and Tumor2. Centralised training on all 5 databases together is also presented.
Setting Mod. Drop Norm. Needs Targ. Data ISLES Tumor2 Average
BN N/A 57.9 75.3 66.6
IN N/A 55.7 76.7 66.2
Upper Bound: Trained on Target only NF N/A 60.6 72.6 66.6
Trained on ATLAS (T1) IN N/A 5.3 -
Trained on BRATS (T1) IN N/A - 63.7 34.5
Centralised IN No 50.7 58.1 54.4
Centralised IN No 55.5 72.2 63.9
Federated (avg. BN params) BN No 53.8 0.0 26.9
Federated (avg. BN params) BN No 43.1 68.0 55.6
Federated IN No 48.3 59.1 53.7
Federated IN No 44.2 72.7 58.5
Federated NF No 54.6 31.1 42.9
Federated NF No 52.8 72.1 62.5
Federated (client spec. BN params) BN Yes 50.2 36.8 43.5
Federated (client spec. BN params) BN Yes 49.9 70.1 60.0
Investigating the Effects of Normalization Layers on Source Clients: Tab. 2 shows the effects of feature normalization on the source databases. The findings indicate that keeping client-specific BN layers yields better performance compared to other normalization techniques on the source clients, but pose difficulties when dealing with unseen clients (as shown in Table 4). The NF method outperforms other non client-specific feature normalization layers in both source client and generalization results. Consequently, a trade-off must be made when designing a model for training with this type of heterogeneous data: either improve the source clients’ performance or prioritizing the development of a model that generalizes to unseen clients.
4 Conclusion
This study shows the feasibility, practicality and effectiveness of using FL to train a single model across decentralised brain MRI databases, each with different input modalities and segmentation tasks. Our results show that federated joint training on multiple databases results in a single model that can segment all diseases and sets of modalities seen in training clients as well as client-specific models trained for a specific set of modalities and disease. Furthermore, it shows promising generalization to unseen clients with modality combinations not seen during training, and highlight the importance of the feature normalization method’s choice to address data heterogeneity. We believe this study highlights the benefits of leveraging heterogeneous brain MRI databases for federated training and opens up future research directions in this area.
5 Acknowledgement
FW is supported by the EPSRC Centre for Doctoral Training in Health Data Science (EP/S02428X/1), by the Anglo-Austrian Society, and by an Oxford-Reuben scholarship. PS is supported in part by the UK EPSRC Programme Grant EP/T028572/1 (VisualAI), in part by the UKRI grant reference EP/X040186/1 (Turing AI Fellowship: Ultra Sound Multi-Modal Video-based Human-Machine Collaboration), and a UK EPSRC Doctoral Training Partnership award. ZL is supported by scholarship provided by the EPSRC Doctoral Training Partnerships programme [EP/W524311/1]. VN, NIHR Rosetrees Trust Advanced Fellowship, NIHR302544, is funded in partnership by the NIHR and Rosetrees Trust. NV is supported by the NIHR Oxford Health Biomedical Research Centre (NIHR203316). The views expressed are those of the authors and not necessarily those of the NIHR, Rosetrees Trust or the Department of Health and Social Care. The authors also acknowledge the use of the University of Oxford Advanced Research Computing (ARC) facility in carrying out this work (http://dx.doi.org/10.5281/zenodo.22558).